### v1

@@ Abstract
The equations for the photon surface in spherical symmetry are worked out, starting from [1], in the most general dynamical setting. We show that the condition for a timelike hypersurface to be a photon surface can be reformulated as a non-autonomous dynamical system, whose analysis reveals that the same condition also holds when the surface is generated by a null radial geodesic.
As an application, we consider a well-known model of a spherical dust cloud undergoing gravitational collapse. Comparing our findings with those in [2], we establish that the photon surface uniquely extends in the interior spacetime as a null hypersurface, allowing us to analytically investigate whether it covers the singularity developing in the LTB model.
@@ 1 Introduction
Detection of black holes represents a intriguing challenge in the realm of observational astrophysics. The well-known international collaboration Event Horizon Telescope (EHT) has utilized a network of radio telescopes across the globe to elaborate an image of the black hole in galaxy M87 [3, 4], often referred to as the black hole shadow. The EHT collaboration has applied analogous methods to investigate other substantial cosmic entities also, notably the SgrA* black hole situated at the Milky Way’s core [5].
@@
It is important to note that the EHT images are not ”direct photographs,” but rather reconstructed radio maps illustrating hot gas in orbit around the black hole, which manifest as a luminous ring.
@@
Indeed, despite its name, what EHT aims to capture is not exactly the shape of the event horizon, but rather something that reveals the motion of objects revolving around the black hole, within the observable region to a faraway observer, which can be effectively described, in the simplest toy models, by the so-called photon sphere.
@@
Classically, the concept of a photon sphere arises in static and spherically symmetric settings such as the Schwarzschild solution, where it manifests as a timelike hypersurface at radius $r=3M$ on which initially tangent null geodesics remain tangent as they evolve. A similar approach, based on the study of null radial geodesics, has been used in a context of static spacetimes [6, 7, 8, 9, 10, 11, 12],
@@
When it comes to dynamical spacetimes, the situation is more involved. One could think to generalize the above approach to this situation, which actually proves fruitful where the metric assumes a special form - for example, [13, 14, 15] study the situation for a spherical metric written in Bondi coordinates $(v,r,\theta,\phi)$, where $v$ is a null coordinate, as is the case for the Vaidya solution.
@@
However, this notion has been extended and generalized to a broader geometric object known as the photon surface, formalized in the seminal work by Claudel, Virbhadra, and Ellis [1], where it is defined independently of any specific symmetry or Killing structure.
@@
In their framework, photon surfaces are understood as nowhere-spacelike hypersurfaces invariant under the flow of null geodesics that remain confined within the surface. This perspective allows for a more flexible analysis, applicable not only in static spacetimes [16] but also in dynamical regimes where symmetry assumptions are weakened or absent. Subsequent studies, such as those by Cao and Song [2], have proposed quasi-local characterizations of photon surfaces in general spherically symmetric spacetimes, aiming to preserve the spirit of Claudel et al.’s geometric definition while adapting it to the complexity of evolving gravitational systems.
@@
One physically rich setting in which photon surfaces can be investigated dynamically is gravitational collapse. The Lemaitre–Tolman–Bondi (LTB) models, describing inhomogeneous spherical dust clouds undergoing collapse under their own gravity, provide a valuable laboratory for probing how such surfaces behave in the presence of singularities and horizons. These models allow for analytical control over the metric and matter profiles and, depending on the initial data, can yield either black hole formation or naked singularities, offering insight into the end-state of light rays and causal curves near regions of extreme curvature [17]. Recent investigations have demonstrated that shadows and photon spheres, traditionally associated with black holes, can also arise in certain naked singularity spacetimes resulting from gravitational collapse, depending on specific parameters of the model [18]. This challenges the conventional interpretation of shadow observations as definitive evidence for black holes.
@@
To carry out this investigation, we start by reviewing the photon surface condition for timelike hypersurfaces in spherically symmetric spacetimes, as given in [1], and derive the relevant differential equations that such surfaces must satisfy. We then apply this framework to the LTB model, examining how the photon surface extends across the boundary between interior and exterior spacetimes and determining the precise conditions under which such an extension exists and remains physically meaningful. In this respect, the present paper aims to compare and discuss its findings with those presented in [2], trying to highlight some aspects that could be otherwise underestimated, particularly regarding the relation between the photon surface behavior and the appearance of naked singularities.
@@ 2 Photon surfaces in spherical symmetry
It is a well-established fact that Schwarzschild spacetime exhibits $r=3m$ as a hypersurface with the property that initially tangent null geodesics remain tangent, called photon sphere. As we shall see, the paper [1] suggests that photon spheres arise as particular cases of more general objects, photon surfaces.
@@
In the spherically symmetric case, [1] states a condition for a timelike hypersurface to be a photon surface:
@@
[1, Theorem III.1] Let $(M,g)$ be a spherically symmetric spacetime and $S$ an SO(3)SO(3)–invariant timelike hypersurface of $(M,g)$ and let $X$ be the SO(3)SO(3)–invariant unit future–directed timelike tangent vector field along $S$ orthogonal to the SO(3)SO(3)–invariant 2–spheres in $S$. Let $\mathcal{T}$ be one such SO(3)SO(3)–invariant 2–sphere in $S$ and $\mathcal{T}_{s}$ be the SO(3)SO(3)–invariant 2–sphere in $S$ at arc length $s$ from $\mathcal{T}$ along the integral curves of $X$. Then $S$ is a photon surface of $(M,g)$ iff
@@
Inspection of (4a)–(4b) above leads us to conclude that, if $y(t_{0})=\pm 1$ for some $t_{0}$ then $y(t)=\pm 1$ throughout the solution. Let us notice that $\dot{x}=\pm e^{\nu-\lambda}$ is the equation of a radial null geodesic in this spacetime. Therefore, although we already know (see Remark 1) that null hypersurfaces are trivially photon surfaces, we can see that equation (1) in Theorem 1 – that in principle assumes the surface to be timelike – also holds for those null hypersurfaces $S$ generated by a radial null geodesic $(t,x(t))$ in the 2–dimensional quotient submanifold $M/S^{2}$. We should remember this fact, which will prove helpful when discussing the initial conditions for equation (3) or, equivalently, the system (4a)–(4b). As we will observe in the examples we explore later, these initial conditions are essentially determined by the boundary conditions applied to the metric being analyzed.
@@
Let us consider the particular case where (2) is static, therefore admitting a $SO(3)\times\mathbb{R}$ group of isometries, where the $\mathbb{R}$ isometries are generated by a Killing vector field $K$, orthogonal to the $SO(3)$ orbits. In this context, the notion of photon sphere is defined in [1] as an $SO(3)\times\mathbb{R}$ invariant photon surface. In the static case, $\nu,\lambda$ and $r$ are functions of $x$ only, and the Killing vector field that generates the isometries of $\mathbb{R}$ is collinear to $\partial_{t}$. Therefore, a photon space is a photon surface with constant $x(t)$, which in view of (4a)–(4b) is implicitly defined by those $x$ satisfying
@@
For instance, in the case of the anisotropic generalizations of de Sitter spacetimes [19] given by  [^1] Please note that in this example we are using $r$ as a coordinate in place of $x$, in compliance with most of the literature on the subject.
@@
equation (5) takes the well-known form (see [1, (67)])
@@ 3 Photon surfaces in dust collapse
Now, let us apply the notions sketched in the above section to the very well-known model of a spherical dust cloud undergoing gravitational collapse. This is modeled by considering a general spherical metric (2) where the energy–momentum tensor is given by $T=-\rho\,\mathrm{d}t\otimes\partial_{t}$ [20, 17]. The metric is given by the Lemaître–Tolman–Bondi (LTB) model
@@
and the collapse evolution is dictated by the PDE
@@
As pointed out in [2, (5.5)], for the LTB model a lengthy but straightforward computation transforms the equation (3) for photon surfaces as
@@
When it comes to assigning initial conditions for this system, we have first to observe that the metric (7) describes only the interior part of the model we are describing, i.e. the dust star. In other words, this metric is defined for $x\in[0,x_{b}]$, where $x_{b}$ represents the boundary of the star. The timelike hypersurface $\Sigma=\{x=x_{b}\}$ must be matched with a suitable exterior solution that describes the matter distribution outside the spherical dust cloud. It is well known that for the exterior region to be Schwarzschild spacetime, the radial pressure, indicated by the $T^{1}_{1}$ component of the energy-momentum tensor, must be zero at $\Sigma$. According to the Einstein field equations, this condition ensures that the Misner–Sharp mass does not depend on time $t$ and remains constant across the matching hypersurface, a necessary condition for matching [21]. Naturally, this applies to the model being analyzed here, where pressures uniformly vanish throughout the spacetime.
From now on we will confine ourselves to the so-called marginally bound case, which amounts to set $k(x)=0$. This choice does not exclude any of the physical implications of the model, as one can demonstrate [22]; however, it allows the equation (8) to explicitly integrate and yield
@@
where also the initial profile of energy density has been fixed in such a way that $r(0,x)=x$. In this way, each timelike hypersurface with fixed $x=x_{0}$ physically represents the shell of dust cloud particles that, at initial time, are located at a proper distance $x_{0}$ from the center of symmetry, and the function $t\mapsto r(t,x_{0})$ will represent the dynamical evolution of this shell as time flows by. The shell collapses under its own gravitational influence and eventually becomes singular at comoving time $t_{s}(x_{0})=\mu(x_{0})^{-1}$, where the energy density $\rho$ diverges, indicating the presence of a spacetime singularity, which can be covered by the presence of the apparent horizon $t_{h}(x_{0})$, described by the equation $r(t_{h}(x),x)=2m(x)$, see Figure 1.
@@
We will refer to this model as the marginally bound LTB collapse. Its endstate, starting from regular initial conditions at $t=0$, is well known in the literature [20]: let us summarize it briefly. Consider a mass profile such that, using (11)
@@
with $a>0$ and $n\in\mathbb{N}\setminus\{0\}$: in this way the initial energy $\rho_{0}(x)$ is regular up to the central shell $x=0$, and decreasing in a right neighborhood of $x=0$ – we can suppose that the star boundary $x_{b}$ belongs to this neighborhood. In this configuration, the non-central singularities, i.e. $t_{s}(x)$ for every $x\in]0,x_{b}]$ are always covered by the apparent horizon. Conversely, the central singularity $t_{s}(0)$ may be either covered or naked, depending on the choice of parameters $a$ and $n$ in (12). Indeed, one has to observe that deriving $t_{h}(x)$ from the condition $r=2m$ leads to the expression
@@
Full details of these studies are based on the qualitative study of radial null geodesics, and can be found in [23, 22]. According to these findings, the central singularity is:
@@
at the star boundary $\Sigma=\{x=x_{b}\}$. Israel–Darmois junction conditions [24] guarantee that the matching is $C^{1}$ across $\Sigma$, provided that the Schwarzschild mass is given by $M=m(x_{b})$. Observe that the coordinates $(\tau,r)$ in (13) will be related to $t$ on the matching surface – in particular, $r=r(t,x_{b})$ on $\Sigma$ [^2] That justifies the slight abuse of notation where we employ the same letter $r$ for both the radial coordinate in (13) and the metric coefficient (11) within (7).. Details can be found in [25].
@@
In view of Remark 1 we can accept $y_{0}=\pm 1$ as initial conditions, since they will correspond to the outgoing ($y_{0}=1$) and ingoing ($y_{0}=-1$) radial null geodesics starting at $x=x_{b}$. Among the above cases, the outgoing curve will necessarily cover the ingoing one, which makes the $y_{0}=1$ extension more significant than the other.
Of course, one may consider other extensions which start spacelike at $x=x_{b}$. These choices would correspond to setting $|y_{0}|<1$. However, $y_{0}<0$ would correspond to a decreasing $x(t)$ curve, and for the same reason would be less significant than the extension that we have considered with $y_{0}=1$. The following lemma will help us rule out all the other cases.
@@
Before presenting the proof, Figure 2 illustrates the core concept it is based on: the (non–autonomous) vector field induced by (10a)–(10b) prevents the existence of curves with initial data $|y_{0}|<1$ that extend back to $x=0$.
@@
After the above intuitive idea, let us start the proof by showing, as an intermediate step, that $\exists t_{1}<t_{0}\,:\,y(t_{1})=0$. By contradiction, suppose that $y(t)>0,\forall t<0$. Then $x(t)$ is increasing $\forall t<t_{0}$ and then, called $t_{\text{inf}}\in\mathbb{R}\cup\{+\infty\}$ the infimum of the maximal interval where the solution can be extended, there exists $\lim_{t\to t_{\text{inf}}}x(t)=x_{0}\in[0,x_{b}[$.
@@
Since the right-hand side positively diverges as $t\to$-$\infty$, we get a contradiction, recalling that $y(t)$ should remain positive. Then $t_{\text{inf}}\in\mathbb{R}$, and this situation takes place only if $\lim_{t\to t_{\text{inf}^{+}}}x(t)=0$. But in this case, again (10a)–(10b) gives
@@
from which one gets the existence of a positive constant $C_{3}$ such that
@@
Then so far we have shown that $y(t_{1})=0$ for some $t_{1}<t_{0}$. Inspection of the equivalent equation (9) shows that if $\dot{x}=0$, then $\ddot{x}>0$, then $x(t_{1})$ is a local minimum for $x(t)$, and so in a left neighborhood of $t_{1}$, $x(t)$ is decreasing and then $y(t)$ is negative. This fact forces $y(t)<0,\,\forall t<t_{1}$ and then $x(t)$ will be decreasing for earlier times than $t_{1}$. Let us show that $x(\bar{t})=x_{b}$ for some $\bar{t}<t_{1}$. If that is not the case, then $(x(t),y(t))$ will remain in an area where the ODE is regular ($x(t)>x(t_{1})$ and $y\in]-1,0]$). This means that the solution can be extended $\forall t<t_{1}$. But in this case the same estimate as (15) gives $y(t)\to-1$, and using (10a) we get the existence of a positive constant $C_{4}$ such that
@@
that integrated as $t\to$-$\infty$ gives $x(t)$ positively diverging, which is a contradiction, and this concludes the proof of the lemma. ∎
The above Lemma shows some extension of photon surfaces that are not physically acceptable. The reason is that these extensions have two points of contact with the star boundary - one at $t_{0}$ and the other one at an earlier comoving time $\bar{t}<t_{0}$. Since we are considering a collapsing model, i.e. $r_{t}<0$, we have $r(\bar{t},x_{b})>r(t_{0},x_{b})=3M$. Therefore, the extension at $\bar{t}$ cannot be matched with the exterior photon space, that indeed is matched at $t=t_{0}$. We conclude that the null extension is the only possible one: the findings of this section can be then summarized in the following.
@@
which is claimed in [2] to be assumed  [^3] Observe that in [2] Misner-Sharp mass is denoted by $E$. as a second initial condition, besides $x(t_{0})=x_{b}$. But actually, that would give $\dot{x}(t_{0})=0$ since the mass is a function of $x$ only and in fact, from the second order ODE (9) we have that, when $\dot{x}(t)=0$ then $\ddot{x}(t)=1/(r\,r_{x})>0$ and so a critical point of the curve $x(t)$ is necessarily a strict local minimum. That excludes this possibility because, if we search for an extension of the photon surface in the interior portion of the model, we would rather look for a function $x(t)<x_{b}$, and this cannot happen in this case where, in a neighborhood of the initial time $t_{0}$, $x(t)$ would attain strictly greater values than $x(t_{0})=x_{b}$.
The apparent contradiction is indeed resolved because, as a matter of fact, the condition used in [2] is precisely that coming from our Theorem 2, as one can verify looking at [2, (5.41)] – although misinterpreted there, as explained above.

### v2

@@ Abstract
The equations for the photon surface in spherical symmetry are worked out, starting from [Claudel:2000yi], in the most general dynamical setting. We show that the condition for a timelike hypersurface to be a photon surface can be reformulated as a non-autonomous dynamical system, whose analysis reveals that the same condition also holds when the surface is generated by a null radial geodesic.
As an application, we consider a well-known model of a spherical dust cloud undergoing gravitational collapse. Comparing our findings with those in [Cao:2019vlu], we establish that the photon surface uniquely extends in the interior spacetime as a null hypersurface, allowing us to analytically investigate whether it covers the singularity developing in the Lemaitre-Tolman-Bondi (LTB) model.
@@ 1 Introduction
Detection of black holes represents an intriguing challenge in the realm of observational astrophysics. The well-known international collaboration Event Horizon Telescope (EHT) has utilized a network of radio telescopes across the globe to elaborate an image of the black hole in galaxy M87 [EventHorizonTelescope:2019dse, EventHorizonTelescope:2024dhe], often referred to as the black hole shadow. The EHT collaboration has also applied analogous methods to investigate other substantial cosmic entities, notably the SgrA black hole situated at the Milky Way’s core [EventHorizonTelescope:2022wkp].
@@
It is important to note that the EHT images are not direct optical photographs but rather reconstructed radio maps that trace the emission from hot gas orbiting the black hole, which gives rise to the observed luminous ring.
@@
Indeed, despite its name, the EHT does not aim to capture the shape of the event horizon itself, but rather to trace the motion of matter orbiting the black hole within the observable region to a faraway observer, which can be effectively described, in the simplest toy models, by the so-called photon sphere.
@@
Classically, the concept of a photon sphere arises in static and spherically symmetric settings such as the Schwarzschild solution, where it manifests as a timelike hypersurface at radius $r=3M$ on which initially tangent null geodesics remain tangent as they evolve. A similar approach, based on the study of null radial geodesics, has been used in the context of static spacetimes [Nolan:2014maa, Hasse:2001by, Chakraborty:2011uj, Bogush:2023ojz, Bogush:2024fqj, Sadeghi:2024itx, Vertogradov:2024dpa], or even in the context of non–transparent collapse [Schneider:2018hge].
@@
When it comes to a fully dynamical setting, the situation is more involved. One could think of generalizing the above approach to this situation, which actually proves fruitful when the metric assumes a special form - for example, [Mishra:2019trb, Solanki:2022glc, Vertogradov:2024eim] study the situation for a spherical metric written in Bondi coordinates $(v,r,\theta,\phi)$, where $v$ is a null coordinate, as is the case for the Vaidya solution.
@@
However, this notion has been extended and generalized to a broader geometric object known as the photon surface, formalized in the seminal work by Claudel, Virbhadra, and Ellis [Claudel:2000yi], where it is defined independently of any specific symmetry or Killing structure.
@@
In their framework, photon surfaces are understood as nowhere-spacelike hypersurfaces invariant under the flow of null geodesics that remain confined within the surface. This perspective allows for a more flexible analysis, applicable not only in static spacetimes [Virbhadra:2002ju] but also in dynamical regimes where symmetry assumptions are weakened or absent. Subsequent studies, such as those by Cao and Song [Cao:2019vlu], have proposed quasi-local characterizations of photon surfaces in general spherically symmetric spacetimes, aiming to preserve the spirit of Claudel et al.’s geometric definition while adapting it to the complexity of evolving gravitational systems.
@@
One physically rich setting in which photon surfaces can be investigated dynamically is gravitational collapse. The Lemaitre–Tolman–Bondi (LTB) models, describing inhomogeneous spherical dust clouds undergoing collapse under their own gravity, provide a valuable laboratory for probing how such surfaces behave in the presence of singularities and horizons. These models allow for analytical control over the metric and matter profiles and, depending on the initial data, can yield either black hole formation or naked singularities, offering insight into the end-state of light rays and causal curves near regions of extreme curvature [Joshi:2008zz]. Recent investigations have demonstrated that shadows and photon spheres, traditionally associated with black holes, can also arise in certain naked singularity spacetimes resulting from gravitational collapse, depending on specific parameters of the model [Shaikh:2018lcc]. This challenges the conventional interpretation of shadow observations as definitive evidence for black holes.
@@
To carry out this investigation, we start by reviewing the photon surface condition for timelike hypersurfaces in spherically symmetric spacetimes, as given in [Claudel:2000yi], and derive the relevant differential equations that such surfaces must satisfy. We then apply this framework to the LTB model, examining how the photon surface extends across the boundary between interior and exterior spacetimes and determining the precise conditions under which such an extension exists and remains physically meaningful. In this respect, the present paper aims to compare and discuss its findings with those presented in [Cao:2019vlu], trying to highlight some aspects that might otherwise be overlooked, particularly regarding the relation between the photon surface behavior and the appearance of naked singularities.
Indeed, recent studies have shown that black holes and naked singularities in this context produce almost indistinguishable shadows [Kong:2013daa, Kong:2013aja]. However, such similarity arises only at late times, whereas the underlying dynamics responsible for the shadows differ between the naked singularity and the black hole configurations [Ortiz:2013wza, Ortiz:2015rma]. In our work, we highlight that the extension of the photon surface behaves differently between black holes and naked singularities, and these differences match the distinct shadow formation dynamics observed in previous research.
@@ 2 Photon surfaces in spherical symmetry
It is a well-established fact that Schwarzschild spacetime exhibits $r=3m$ as a hypersurface with the property that initially tangent null geodesics remain tangent, known as photon sphere. As we shall see, the paper [Claudel:2000yi] suggests that photon spheres arise as particular cases of more general objects, photon surfaces.
@@
In the spherically symmetric case, [Claudel:2000yi] states a condition for a timelike hypersurface to be a photon surface:
@@
[Claudel:2000yi, Theorem III.1] Let $(M,g)$ be a spherically symmetric spacetime and $S$ an SO(3)SO(3)–invariant timelike hypersurface of $(M,g)$, $X$ be the SO(3)SO(3)–invariant unit future–directed timelike tangent vector field along $S$ orthogonal to the SO(3)SO(3)–invariant 2–spheres in $S$. Let $\mathcal{T}$ be one such SO(3)SO(3)–invariant 2–sphere in $S$ and $\mathcal{T}_{s}$ be the SO(3)SO(3)–invariant 2–sphere in $S$ at arc length $s$ from $\mathcal{T}$ along the integral curves of $X$. Then $S$ is a photon surface of $(M,g)$ iff
@@
Inspection of (4a)–(4b) above leads us to conclude that if $y(t_{0})=\pm 1$ for some $t_{0}$, then $y(t)=\pm 1$ throughout the solution. Let us notice that $\dot{x}=\pm e^{\nu-\lambda}$ is the equation of a radial null geodesic in this spacetime. Therefore, although we already know (see Remark 1) that null hypersurfaces are trivially photon surfaces, we can see that equation (1) in Theorem 1, which in principle assumes the surface to be timelike, also holds for those null hypersurfaces $S$ generated by a radial null geodesic $(t,x(t))$ in the 2–dimensional quotient submanifold $M/S^{2}$. We should remember this fact, which will prove helpful when discussing the initial conditions for equation (3) or, equivalently, the system (4a)–(4b). As we will observe in the examples we explore later, these initial conditions are essentially determined by the boundary conditions applied to the metric being analyzed.
@@
Let us consider the particular case where (2) is static, therefore admitting a $SO(3)\times\mathbb{R}$ group of isometries, where the $\mathbb{R}$ isometries are generated by a Killing vector field $K$, orthogonal to the $SO(3)$ orbits. In this context, the notion of photon sphere is defined in [Claudel:2000yi] as an $SO(3)\times\mathbb{R}$ invariant photon surface. In the static case, $\nu,\lambda$ and $r$ are functions of $x$ only, and the Killing vector field that generates the isometries of $\mathbb{R}$ is collinear to $\partial_{t}$. Therefore, a photon space is a photon surface with constant $x(t)$, which in view of (4a)–(4b) is implicitly defined by those $x$ satisfying
@@
For instance, in the case of the anisotropic generalizations of de Sitter spacetimes [Giambo:2002wr] given by  [^1] Please note that in this example we are using $r$ as a coordinate in place of $x$, in compliance with most of the literature on the subject.
@@
equation (5) takes the well-known form (see [Claudel:2000yi, (67)])
@@ 3 Photon surfaces in dust collapse
Now, let us apply the notions sketched in the previous section to the well-known model of a spherical dust cloud undergoing gravitational collapse. This is obtained by considering a general spherical metric (2) where the energy–momentum tensor is given by $T=-\rho\,\mathrm{d}t\otimes\partial_{t}$ [Singh:1994tb, Joshi:2008zz]. The metric is given by the Lemaître–Tolman–Bondi (LTB) model
@@
where the collapse evolution is dictated by the PDE
@@
As pointed out in [Cao:2019vlu, (5.5)], for the LTB model a lengthy but straightforward computation transforms the equation (3) for photon surfaces as
@@
When it comes to assigning initial conditions for this system, we first have to observe that the metric (7) describes only the interior region of the model, i.e., the dust star. In other words, this metric is defined for $x\in[0,x_{b}]$, where $x_{b}$ represents the boundary of the star. The timelike hypersurface $\Sigma=\{x=x_{b}\}$ must be matched with a suitable exterior solution that describes the matter distribution outside the spherical dust cloud. It is well known that for the exterior region to be Schwarzschild spacetime, the radial pressure, indicated by the $T^{1}_{1}$ component of the energy-momentum tensor, must be zero at $\Sigma$. According to the Einstein field equations, this condition ensures that the Misner–Sharp mass does not depend on time $t$ and remains constant across the matching hypersurface, a necessary condition for matching [Magli_1997]. Naturally, this applies to the model being analyzed here, where pressures uniformly vanish throughout the spacetime.
From now on, we will confine ourselves to the so-called marginally bound case, which amounts to setting $k(x)=0$. On the one hand, this choice does not exclude any of the physical implications of the model because, as one can demonstrate [Giambo:2002tp], the spacetime causal structure is reduced to the study of a dynamical system depending on the choice of some parameters typically related to the Taylor expansion of the mass function. This is an instance where a symmetry‑reduced Einstein system exhibits qualitatively distinct regimes separated in parameter space; see [Cotsakis:2023emv, Cotsakis2025]. On the other hand, however, the marginally bound assumption allows the equation (8) to integrate explicitly and yield
@@
where the initial energy-density profile has also been chosen so that $r(0,x)=x$. In this way, each timelike hypersurface with fixed $x=x_{0}$ physically represents the shell of dust cloud particles that, at the initial time, are located at a proper distance $x_{0}$ from the center of symmetry, and the function $t\mapsto r(t,x_{0})$ will represent the dynamical evolution of this shell as time flows by. The shell collapses under its own gravitational influence and eventually becomes singular at comoving time $t_{s}(x_{0})=\mu(x_{0})^{-1}$, where the energy density $\rho$ diverges, indicating the presence of a spacetime singularity, which can be covered by the presence of the apparent horizon $t_{h}(x_{0})$, described by the equation $r(t_{h}(x),x)=2m(x)$, see Figure 1.
@@
We will refer to this model as the marginally bound LTB collapse. Its end state, starting from regular initial conditions at $t=0$, is well known in the literature [Singh:1994tb]: let us summarize it briefly. Consider a mass profile such that, using (11)
@@
with $a>0$ and $n\in\mathbb{N}\setminus\{0\}$, so the initial energy $\rho_{0}(x)$ is regular up to the central shell $x=0$, and decreasing in a right neighborhood of $x=0$ – we can suppose that the star boundary $x_{b}$ belongs to this neighborhood. In this configuration, the non-central singularities, i.e., $t_{s}(x)$ for every $x\in]0,x_{b}]$ are always covered by the apparent horizon. Conversely, the central singularity $t_{s}(0)$ may be either covered or naked, depending on the choice of parameters $a$ and $n$ in (12). Indeed, one has to observe that deriving $t_{h}(x)$ from the condition $r=2m$ leads to the expression
@@
Full details of these studies are based on the qualitative study of radial null geodesics and can be found in [Mena:2001dr, Giambo:2002tp]. According to these findings, the central singularity is:
@@
at the star boundary $\Sigma=\{x=x_{b}\}$. Israel–Darmois junction conditions [Israel:1966rt] is sufficient to ensure that the matching is $C^{1}$ across $\Sigma$, provided that the Schwarzschild mass is given by $M=m(x_{b})$. Observe that the coordinates $(\tau,r)$ in (13) will be related to $t$ on the matching surface – in particular, $r=r(t,x_{b})$ on $\Sigma$ [^2] That justifies the slight abuse of notation where we employ the same letter $r$ for both the radial coordinate in (13) and the metric coefficient (11) within (7).. Details can be found in [Giambo2005GravitationalFields].
@@
In view of Remark 1, we can accept $y_{0}=\pm 1$ as initial conditions, since they will correspond to the outgoing ($y_{0}=1$) and ingoing ($y_{0}=-1$) radial null geodesics starting at $x=x_{b}$. Among the above cases, the outgoing curve will necessarily cover the ingoing one, which makes the $y_{0}=1$ extension more significant than the $y_{0}=-1$ one.
Of course, one may consider other extensions that start spacelike at $x=x_{b}$. These choices correspond to setting $|y_{0}|<1$. However, $y_{0}<0$ would correspond to a decreasing $x(t)$ curve and would, for the same reason, be less significant than the $y_{0}=1$ extension. The following lemma will help us rule out all the other cases.
@@
Before presenting the proof, we illustrate its core concept in Figure 2: the (non–autonomous) vector field induced by (10a)–(10b) prevents the existence of curves with initial data $|y_{0}|<1$ that extend back to $x=0$.
@@
After the above intuitive idea, let us start the proof by showing, as an intermediate step, that $\exists t_{1}<t_{0}\,:\,y(t_{1})=0$. By contradiction, suppose that $y(t)>0,\forall t<0$. Then $x(t)$ is increasing $\forall t<t_{0}$ and then, called $t_{\text{inf}}\in\mathbb{R}\cup\{-\infty\}$ the infimum of the maximal interval where the solution can be extended, there exists $\lim_{t\to t_{\text{inf}}^{+}}x(t)=x_{0}\in[0,x_{b}[$.
@@
Since the right-hand side positively diverges as $t\to$-$\infty$, we get a contradiction, recalling that $y(t)$ should remain positive. Then $t_{\text{inf}}\in\mathbb{R}$, and this situation may occur only when there is a loss of regularity in the system, that is, if $\lim_{t\to t_{\text{inf}^{+}}}x(t)=0$. But in this case, again (10a)–(10b) gives
@@
from which one has the existence of a positive constant $C_{3}$ such that
@@
Then so far we have shown that $y(t_{1})=0$ for some $t_{1}<t_{0}$. Inspection of the equivalent equation (9) shows that, if $\dot{x}=0$, then $\ddot{x}>0$, so $x(t_{1})$ is a local minimum for $x(t)$. Hence, in a left neighborhood of $t_{1}$, $x(t)$ is (strictly) decreasing and therefore $y(t)$ is negative. This fact forces $y(t)<0,\,\forall t<t_{1}$ and then $x(t)$ will be decreasing for earlier times than $t_{1}$. Let us show that $x(\bar{t})=x_{b}$ for some $\bar{t}<t_{1}$. If that is not the case, then $(x(t),y(t))$ will remain in an area where the ODE is regular ($x(t)>x(t_{1})$ and $y\in]-1,0]$). This means that the solution can be extended $\forall t<t_{1}$. But in this case the same estimate as (15) gives $y(t)\to-1$, and using (10a) we get the existence of a positive constant $C_{4}$ such that
@@
that integrated as $t\to$-$\infty$ gives $x(t)$ positively diverging, which is a contradiction. This concludes the proof of the lemma. ∎
The above Lemma shows extensions of photon surfaces that are not physically acceptable. The reason is that these extensions have two points of contact with the star boundary – one at $t_{0}$ and the other at an earlier comoving time $\bar{t}<t_{0}$. Since we are considering a collapsing model, i.e. $r_{t}<0$, we have $r(\bar{t},x_{b})>r(t_{0},x_{b})=3M$. Therefore, the extension at $\bar{t}$ cannot be matched with the exterior photon space, which is indeed matched at $t=t_{0}$. We conclude that the null extension is the only possible one; the findings of this section can then be summarized as follows.
@@
which is claimed in [Cao:2019vlu] to be assumed  [^3] Observe that in [Cao:2019vlu] the Misner-Sharp mass is denoted by $E$. as a second initial condition, in addition to $x(t_{0})=x_{b}$. But actually, that would give $\dot{x}(t_{0})=0$ since the mass depends only on $x$; moreover from the second order ODE (9) we have that when $\dot{x}(t)=0$ then $\ddot{x}(t)=1/(r\,r_{x})>0$ and so a critical point of the curve $x(t)$ is necessarily a strict local minimum. That excludes this possibility because, if we search for an extension of the photon surface in the interior portion of the model, we would rather look for a function $x(t)<x_{b}$, and this cannot happen in this case where, in a neighborhood of the initial time $t_{0}$, $x(t)$ would attain strictly greater values than $x(t_{0})=x_{b}$.
The apparent contradiction is indeed resolved because, as a matter of fact, the condition used in [Cao:2019vlu] is precisely the one derived from our Theorem 2, as one can verify from [Cao:2019vlu, (5.41)], although it is misinterpreted there, as explained above.